The price of a 2-hour plumbing repair job is $140 and the price of a 5-hour job is $290. Use the slope to predict the cost of a repair job lasting 4 hours.
Answer by stanbon(48558)

The price of a 2-hour plumbing repair job is $140 and the price of a 5-hour job is $290. Use the slope to predict the cost of a repair job lasting 4 hours.
--------------------
You have two points relating hours and cost:
(2,140) and (5,290)
--------------
slope = (290-140)/(5-2) = 150/3 = 50
Form: y = mx + b
140 = 50*2 + b
b = 40
----------------
EQUATION:
Cost(x) = 50x + 40
C(4) = 5*4 + 40
C(4) = $60
===============
Cheers,
Stan H.